Circular No. 7479

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 O3 (SOHO) Further to IAUC 7472, D. Hammer has provided measurements of a

comet detected by the SOHO C2 and C3 instruments and found by J. D. Shanklin via the SOHO website. The reduced measurements and orbits by B. G. Marsden, together with a search ephemeris, are given on MPEC 2000-Q09. G. J. Garradd, Loomberah, N.S.W., reports that his search for this object around Aug. 21.4 UT, out to ∼ 0^{o}.5 ahead of its predicted position, yielding nothing to mag ∼ 18.

2000 UT  α_{2000}  δ_{2000}
July 30.221  8^{h}21.^{m}5  +19^{o}08′

COMETS C/2000 P1 AND C/2000 P2 (SOHO) Hammer also reports the measurements of two more Kreutz sungraz-

ing comets found by M. Boschat in SOHO C3 website images, with the discovery positions given below. MPEC 2000-Q08 contains the reductions and orbital elements by Marsden.

Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 P1 Aug. 6.654  8^{h}52.^{m}0  +15^{o}16′
C/2000 P2  14.721  9 19.2  +12 49

SATELLITES OF URANUS Following the recoveries of the 1999 satellites of Uranus (cf. IAUC

7447, 7450, 7473), permanent designations (and names) were approved at the IAU General Assembly in Manchester, England: Uranus XVIII = S/1999 U 3 (Prospero), Uranus XIX = S/1999 U 1 (Setebos), Uranus XX = S/1999 U 2 (Stephano).

EDITORIAL NOTICE At the meeting of IAU Commission 6 in Manchester on Aug. 15, Brian

G. Marsden announced his retirement, effective the next day, as Director of the Central Bureau. His place was taken by the undersigned. Marsden had been in charge of the Bureau since Jan. 1968. In recognition of his long service, he has been given the honorary title of Director Emeritus, and it is anticipated that he will still be actively involved in the Bureau’s activities. He has also been elected President of Commission 6.

2000 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green